OpenStudy (anonymous):

Write each expression in the standard form for a complex number, a + bi. [(1 / 2)(cos(72°)) + isin(72°)]5 [√1(cos(3π /14)) + isin(3π /14)]7 [√5(cos(5π / 16)) + isin(5π / 16)]4 [3√7(cos(π / 18)) + isin(π / 18)]6

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@geerky42 Can you help me? You don't have to explain every problem, just help me understand the basics

OpenStudy (anonymous):

take the number out front and raise it to the power then multiply the angle by the power

OpenStudy (anonymous):

Seems simple enough, is that all?

OpenStudy (anonymous):

\[(\frac{1}{2}\left(\cos(72^\circ)+i\sin(72^\circ)\right)^5\] \[=\frac{1}{2^5}\left(\cos(72\times 5)+i\sin(72\times 5)\right)\]

OpenStudy (anonymous):

yes it is real easy of course you still have to do the arithmetic, then evaluate the functions

OpenStudy (anonymous):

\[\frac{1}{32}\left(\cos(360)+i\sin(360)\right)\] \[\frac{1}{32}\left(1+0\right)\]\[\frac{1}{32}\]

OpenStudy (anonymous):

Seems simple enough, is that all?